OpenStudy (christos):

How can I find the domain of the function g(x) = x / x^2 - 1 ?

OpenStudy (christos):

without a graph

terenzreignz (terenzreignz):

Usually, in domain problems, you get expressions in radicals or in denominators. In your case, you have a denominator x^2 - 1 In these cases, the domain is all real numbers EXCEPT those that would make your denominator zero. When is your denominator zero?

OpenStudy (christos):

when is 1

OpenStudy (christos):

the value of x

terenzreignz (terenzreignz):

Yes... and there's another. Indeed, if x = 1 x^2 - 1 = 0 What else?

OpenStudy (christos):

and -1 ? :D

terenzreignz (terenzreignz):

That's right :) And those are the only ones... x^2 - 1 = (x+1)(x-1) = 0 So, your domain is all real numbers EXCEPT those that would make your denominator zero, which are 1 and -1 Good job :D

OpenStudy (christos):

can it be expressed as [X|X <> -1,1] ?

terenzreignz (terenzreignz):

I prefer R - {-1 , 1} No, it can't be expressed like that because numbers BETWEEN -1 and 1 are still in the domain.

OpenStudy (christos):

hmm here the answer is demostrated as {X|X <> -1,1} but I dont know why

terenzreignz (terenzreignz):

For one thing, I don't know what that notation means :(

OpenStudy (phi):

I think <> is another way to say ≠

OpenStudy (christos):

yes its ≠

terenzreignz (terenzreignz):

Well, if that's the case, then go for it, it's correct :)

OpenStudy (christos):

but why ? :D you said the numbers between are still in the domain?

terenzreignz (terenzreignz):

Yeah, they are, but if what you say is true, that <> is the same as writing ≠ then the set {X|X <> -1,1} is correct because it is read as "the set of all x such that x is not equal to -1 or 1" Which does not exclude numbers between -1 and 1 :)

OpenStudy (christos):

Would it be easy for you to explain to me the meaning of | between those 2 x ? It's the first time I am seeing such a notation

terenzreignz (terenzreignz):

Okay, that's what you call the rule method of listing sets, I think... Takes this form: {x | x <bla bla bla>} I'll take you through it part-by-part {x (the set of all x) | (such that) x <bla bla bla> } (x fulfills some sort of condition) For example, the set of all even integers may be denoted as {x | x is an integer, x = 2k, for some integer k}

terenzreignz (terenzreignz):

By the way, it need not be just one parameter (x). For instance, the set of all complex numbers may be denoted as {a+bi | a and b are real numbers, i is the square root of negative 1}

OpenStudy (christos):

Can you help me translate {X|X ≠ -1,1} into words?